OpenStudy (anonymous):

solving equations with variables on both sides how do u solve 2r-(5-r)=13+2r

OpenStudy (anonymous):

The way to begin is to "collect" terms together... get the "r" terms together and the number terms together: 2r - (5-r) = 13 + 2r 2r -5 + r = 13 + 2r 3r - 5 = 13 + 2r So now the original expression is simplified somewhat. Next, you need to get all the r terms on one side and the numbers on the other.

OpenStudy (anonymous):

do u divide